### v2

@@ Abstract
keywords:
Randomized iterative method, least-squares problem, randomized Kaczmarz
{AMS}
65F10, 65C05
@@ 1 Introduction
(1) $\min_{\boldsymbol{x}\in{}^{d}}\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}\quad\text{for }\boldsymbol{A}\in{}^{n\times d}\text{ and }\boldsymbol{b}\in{}^{n}\text{ with }n>d$
@@
Row-access methods have been proposed as a practical way to solve large least-squares problems. These methods access and process one or a few rows of $\boldsymbol{A}$ at a time. An example of a row-access method is randomized Kaczmarz (RK) [26], which is reviewed in Section 1.1. RK converges exponentially fast if the least-squares problem is consistent, $\boldsymbol{b}=\boldsymbol{A}\boldsymbol{x}_{\star}$ [26, 16]. However, in the inconsistent case $\boldsymbol{b}\neq\boldsymbol{A}\boldsymbol{x}_{\star}$, RK only converges up to a finite horizon. This paper overcomes the finite horizon by combining RK with tail averaging, resulting in a new tail-averaged randomized Kaczmarz (TARK) method.
@@ 1.1 Randomized Kaczmarz
    (2a) $\operatorname{\mathbb{P}}\{i_{t}=i\}=\frac{\left\|\boldsymbol{a}_{i}\right\|^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}}\quad\text{for }i=1,\ldots,n.$
@@
    (2b) $\boldsymbol{x}_{t+1}\coloneqq\boldsymbol{x}_{t}+\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{t}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}}.$
@@
RK can be interpreted as an optimized version of stochastic gradient descent for linear least-squares problems that uses nonuniform selection probabilities to improve the convergence rate and eliminate the need for step size tuning [20]. These probabilities can be precomputed using a single pass through the matrix $\boldsymbol{A}$, which might be expensive. Sometimes this initial computation can be avoided by using rejection sampling [20, Sec. 3]. Alternatively, RK can be implemented with uniform sampling, which is equivalent to applying RK to the diagonally reweighted least-squares problem $\min_{\boldsymbol{x}}\left\|\boldsymbol{D}\boldsymbol{b}-(\boldsymbol{D}\boldsymbol{A})\boldsymbol{x}\right\|^{2}$ for $\boldsymbol{D}=\operatorname{diag}(1/\left\|\boldsymbol{a}_{i}\right\|)$.
@@
Theorem 1.1 (Randomized Kaczmarz: Convergence to a horizon [28]).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration Eq. 2 converges exponentially fast until reaching a finite horizon related to the inconsistency:
$\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\right\|^{2}\leq\underbrace{\left(1-\kappa_{\rm dem}^{-2}\right)^{t}\cdot\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}}_{\text{exponential convergence}}+\underbrace{\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}}_{\text{finite horizon}}.$
Unfortunately, the finite convergence horizon cannot be eliminated without changing the RK algorithm. To overcome this obstacle, several variants of the RK method have been proposed:
  • Randomized extended Kaczmarz [28] accesses and manipulates the columns of $\boldsymbol{A}$ to achieve exponential convergence to $\boldsymbol{x}_{\star}$, even in the inconsistent case. Yet, the column manipulations are prohibitively expensive for the largest problems.
@@
The limitations of these existing methods will be further demonstrated through the experiments in Section 2.3.
@@ 1.2 Tail-averaged randomized Kaczmarz
(3) $\boldsymbol{\overline{x}}_{t}\coloneqq\frac{1}{t-t_{\rm b}}\sum\nolimits_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s},$
which depends on the burn-in time $t_{\rm b}$ and the final time $t$. Tail averaging is frequently applied in Markov chain Monte Carlo [17] to obtain a convergent estimator from stochastically varying samples. Tail averaging has also been combined with numerical optimization methods [3, Thm. 3.2], and it leads to the optimal $\mathcal{O}(1/t)$ convergence rate for stochastic gradient descent for strongly convex loss functions [22, 24, 14].
@@
4:     $\boldsymbol{x}_{s+1}=\boldsymbol{x}_{s}+\frac{b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s}}{\left\|\boldsymbol{a}_{i}\right\|^{2}}\boldsymbol{a}_{i}$
@@
6:$\boldsymbol{\overline{x}}_{t}=\frac{1}{t-t_{\rm b}}\sum_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}$
7:return $\boldsymbol{\overline{x}}_{t}$
Our main proposal is tail-averaged randomized Kaczmarz (TARK), which outputs the tail average Eq. 3 of the standard RK iterates Eq. 2; see Algorithm 1. A variant of TARK for ridge regression problems will be presented in Section 3.
@@
Theorem 1.2 (Mean square error bound for TARK).
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\underbrace{\left(1-\kappa_{\rm dem}^{-2}\right)^{t_{\rm b}}\cdot\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}}_{\text{exponential convergence}}+\underbrace{\frac{2\kappa_{\rm dem}^{2}-1}{t-t_{\rm b}}\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}}_{\text{polynomial convergence}}.$
The proof of Theorem 1.2 appears in Section 2.1.
Similar to MCMC error bounds, Theorem 1.2 decomposes the mean square error into the sum of a bias term that decays exponentially in the burn-in time $t_{\rm b}$ and a variance term that decays as $1/t$ in the final time $t$. In particular, TARK converges when both $t_{\rm b}$ and $t-t_{\rm b}$ go to infinity. To control both terms in this error bound, we recommend selecting $t_{\rm b}\in[t/4,t/2]$; see Appendix A for a storage-efficient implementation that ensures this condition when the final time $t$ is not known in advance.
@@
A similar proof guarantees that TARK converges when $t$ goes to infinity with $t_{\rm b}$ fixed. We have the following alternative version of Theorem 1.2:
Theorem 1.3 (Alternative TARK error bound).
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\frac{2\kappa_{\rm dem}^{2}-1}{t-t_{\rm b}}\cdot\biggl{[}\frac{\left(1-\kappa_{\rm dem}^{-2}\right)^{t_{\rm b}}}{\kappa_{\rm dem}^{2}(t-t_{\rm b})}\cdot\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}+\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}\biggr{]}.$
The proof of Theorem 1.3 appears in Section 2.1.
@@ 2 Analysis and evaluation of tail-averaged randomized Kaczmarz
This section provides a more detailed discussion of TARK. Section 2.1 proves Theorem 1.2, Section 2.2 discusses the optimal convergence rate for row-access methods, Section 2.3 provides numerical experiments, and Section 2.4 extends TARK to semi-infinite least-squares problems.
@@ 2.1 Proof of main theorem
The proof of Theorem 1.2 follows the pattern of analysis initiated in [26], but it takes a step further by bounding the inner product terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{t+s}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{t}-\boldsymbol{x}_{\star})\bigr{]}$ which decay exponentially fast with $s$. For ease of reading, the analysis is presented as three lemmas followed by one main calculation.
Lemma 2.1 (Multi-step expectations).
The RK iteration Eq. 2 satisfies
@@
Proof 2.2.
For any $t\in\{r,\ldots,s-1\}$, write the one-step update Eq. 2b as
@@
Use the sampling probabilities Eq. 2a to calculate the expectation over the random index $i_{t}$:
@@
Iterating this equation completes the proof.
Lemma 2.3 (Demmel condition number bound).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration Eq. 2 satisfies
@@
Proof 2.4.
By the construction of the RK iterates Eq. 2b, observe that $\boldsymbol{x}_{r}$ is in the range of $\boldsymbol{A}^{\top}$, as is the solution vector $\boldsymbol{x}_{\star}=\boldsymbol{A}^{+}\boldsymbol{b}=\boldsymbol{A}^{\top}(\boldsymbol{A}\boldsymbol{A}^{\top})^{+}\boldsymbol{b}$. Hence, $\boldsymbol{x}_{r}-\boldsymbol{x}_{\star}\in\operatorname{range}(\boldsymbol{A}^{\top})$. The result follows by expanding $\boldsymbol{x}_{r}-\boldsymbol{x}_{\star}$ in the basis of $\boldsymbol{A}$’s right singular vectors.
Lemma 2.5 (Mean square errors, modified from [26]).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration Eq. 2 satisfies
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{x}_{r}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{r}\cdot\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}+\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$
@@
Proof 2.6.
For any $t\in\{0,1,\ldots,r-1\}$, write the one-step update Eq. 2b as
(4) $\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\star}=\underbrace{\biggl{[}\mathbf{I}-\frac{\boldsymbol{a}_{i_{t}}^{\vphantom{\top}}\boldsymbol{a}_{i_{t}}^{\top}}{\lVert\boldsymbol{a}_{i_{t}}\rVert^{2}}\biggr{]}}_{\text{orthogonal projection}}\bigl{(}\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\bigr{)}+\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{\star}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}}.$
The decomposition explicitly identifies an orthogonal projection matrix. The matrix is idempotent, it annihilates the vector $\boldsymbol{a}_{i_{t}}$, and it preserves all vectors orthogonal to $\boldsymbol{a}_{i_{t}}$. Hence, using the orthogonal decomposition Eq. 4 it follows
@@
Use the sampling probabilities Eq. 2a to calculate the expectation over the random index $i_{t}$:
@@
$\leq(1-\kappa_{\rm dem}^{-2})\cdot\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\rVert^{2}+\frac{\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}},$
where the inequality follows from Lemma 2.3. Next, take the expectation over the random indices $i_{0},\ldots,i_{t}$:
$\operatorname{\mathbb{E}}\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\star}\rVert^{2}\leq(1-\kappa_{\rm dem}^{-2})\cdot\operatorname{\mathbb{E}}\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\rVert^{2}+\frac{\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}},\quad\text{for each }t\in\{0,\ldots,r-1\}.$
Since $\sum_{s=0}^{\infty}(1-\kappa_{\rm dem}^{-2})^{s}=\kappa_{\rm dem}^{2}=\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$, this equation implies the desired result.
Proof 2.7 (Proof of Theorem 1.2 and Theorem 1.3).
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}=\frac{1}{(t-t_{\rm b})^{2}}\sum_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\star})\bigr{]}.$
Next analyze the terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\star})\bigr{]}$ for $r\leq s$ using Lemmas 2.1, 2.3, and 2.5:
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}$ $=\frac{1}{(t-t_{\rm b})^{2}}\sum_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\star})\bigr{]}$
$\leq\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{t_{\rm b}}\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}+\frac{\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}}{(t-t_{\rm b})^{2}}\sum_{r,s=t_{\rm b}}^{t-1}\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{|s-r|}$
@@
which completes the proof of Theorem 1.2. Theorem 1.3 is proved in a similar way, by explicitly averaging over term A also.

### v3

@@ Abstract
journal: Nuclear Physics B
[…]
[label1]organization=Department of Computing and Mathematical Sciences, California Institute of Technology,addressline=1200 E California Blvd, city=Pasadena, postcode=91125, state=CA, country=USA
[…]
[label2]organization=Department of Mathematics, University of California Berkeley,addressline=110 Sproul Hall, city=Berkeley, postcode=94720, state=CA, country=USA
[…]
[label3]organization=Department of Mathematics, University of California San Diego,addressline=9500 Gilman Drive, city=La Jolla, postcode=92093, state=CA, country=USA
@@ 1 Introduction
$\min_{\boldsymbol{x}\in{}^{d}}\,\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}\quad\text{for }\boldsymbol{A}\in{}^{n\times d}\text{ and }\boldsymbol{b}\in{}^{n}\text{ with }n>d$ (1)
@@
Row-access methods have been proposed as a practical way to solve large least-squares problems. These methods access and process one or a few rows of $\boldsymbol{A}$ at a time. An example of a row-access method is randomized Kaczmarz (RK) [1], which is reviewed in Subsection 1.1. RK converges exponentially fast if the least-squares problem is consistent, $\boldsymbol{b}=\boldsymbol{A}\boldsymbol{x}_{\star}$ [1, 2]. However, in the inconsistent case where $\boldsymbol{b}\neq\boldsymbol{A}\boldsymbol{x}_{\star}$, RK only converges up to a finite horizon. This paper overcomes the finite horizon by combining RK with tail averaging, resulting in a new tail-averaged randomized Kaczmarz (TARK) method.
@@ 1.1 Randomized Kaczmarz
    $\operatorname{\mathbb{P}}\{i_{t}=i\}=\frac{\left\|\boldsymbol{a}_{i}\right\|^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}}\quad\text{for }i=1,\ldots,n.$ (2a)
@@
    $\boldsymbol{x}_{t+1}\coloneqq\boldsymbol{x}_{t}+\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{t}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}}.$ (2b)
@@
RK can be interpreted as an optimized version of stochastic gradient descent for linear least-squares problems that uses nonuniform selection probabilities to improve the convergence rate and eliminate the need for step size tuning [3]. These probabilities can be precomputed using a single pass through the matrix $\boldsymbol{A}$, which might be expensive. Sometimes this initial computation can be avoided by using rejection sampling [3, Sec. 3]. Alternatively, RK can be implemented with uniform sampling, which is equivalent to applying RK to the diagonally reweighted least-squares problem $\min_{\boldsymbol{x}}\,\left\|\boldsymbol{D}\boldsymbol{b}-(\boldsymbol{D}\boldsymbol{A})\boldsymbol{x}\right\|^{2}$ for $\boldsymbol{D}=\operatorname{diag}(1/\left\|\boldsymbol{a}_{i}\right\|)$.
@@
Theorem 1 (Randomized Kaczmarz: Convergence to a horizon [4]).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration (2) converges exponentially fast until reaching a finite horizon related to the inconsistency:
$\operatorname{\mathbb{E}}\,\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\right\|^{2}\leq\underbrace{\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{t}\,\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}}_{\text{exponential convergence}}+\underbrace{\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}}_{\text{finite horizon}}.$
Unfortunately, the finite convergence horizon cannot be eliminated without changing the RK algorithm. To overcome this obstacle, several variants of RK have been proposed:
  • Randomized extended Kaczmarz [4] manipulates the columns of $\boldsymbol{A}$ to achieve exponential convergence to $\boldsymbol{x}_{\star}$, even in the inconsistent case. Yet the column manipulations are prohibitively expensive for the largest problems.
@@
The limitations of these existing methods will be demonstrated through the experiments in Subsection 2.3.
@@ 1.2 Tail-averaged randomized Kaczmarz
$\overline{\boldsymbol{x}}_{t}\coloneqq\frac{1}{t-t_{\rm b}}\sum\nolimits_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s},$ (3)
which depends on the burn-in time $t_{\rm b}$ and the final time $t$. Tail averaging is frequently applied in Markov chain Monte Carlo [10] to obtain a convergent estimator from stochastically varying samples. Tail averaging has also been combined with numerical optimization methods [11, Thm. 3.2], and it leads to the optimal $\mathcal{O}(1/t)$ convergence rate for stochastic gradient descent for strongly convex loss functions [12, 13, 14].
@@
4:     $\boldsymbol{x}_{s+1}=\boldsymbol{x}_{s}+(b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s})\,\boldsymbol{a}_{i}/\left\|\boldsymbol{a}_{i}\right\|^{2}$
@@
6:$\overline{\boldsymbol{x}}_{t}=\bigl{(}\sum_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}\bigr{)}/(t-t_{b})$
7:return $\overline{\boldsymbol{x}}_{t}$
Our main proposal is tail-averaged randomized Kaczmarz (TARK), which outputs the tail average (3) of the standard RK iterates (2); see Algorithm 1. A variant of TARK for ridge regression problems will be presented in Subsection 3.
@@
Theorem 2 (Mean square error bound for TARK).
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\underbrace{\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{t_{\rm b}}\,\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}}_{\text{exponential convergence}}+\underbrace{\frac{2\kappa_{\rm dem}^{2}-1}{t-t_{\rm b}}\,\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}}_{\text{polynomial convergence}}.$
The proof of Theorem 2 appears in Subsection 2.1.
Similar to MCMC error bounds, Theorem 2 decomposes the mean square error into the sum of a bias term that decays exponentially in the burn-in time $t_{\rm b}$ and a variance term that decays as $1/t$ in the final time $t$. In particular, TARK converges when both $t_{\rm b}$ and $t-t_{\rm b}$ go to infinity. To control both terms in this error bound, we recommend selecting $t_{\rm b}\in[t/4,t/2]$; see A for a storage-efficient implementation that ensures this condition when the final time $t$ is not known in advance.
@@
A similar proof guarantees that TARK converges when $t$ goes to infinity with $t_{\rm b}$ fixed. We have the following alternative version of Theorem 2:
Theorem 3 (Alternative TARK error bound).
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\frac{2\kappa_{\rm dem}^{2}-1}{t-t_{\rm b}}\biggl{[}\frac{\kappa_{\rm dem}^{2}\left(1-\kappa_{\rm dem}^{-2}\right)^{t_{\rm b}}}{(t-t_{\rm b})}\,\left\|\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\right\|^{2}+\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}\biggr{]}.$
The proof of Theorem 3 appears in Subsection 2.1.
@@ 2 Analysis and evaluation of tail-averaged randomized Kaczmarz
This section provides a more detailed discussion of TARK. Subsection 2.1 proves Theorems 2 and 3, Subsection 2.2 discusses the optimal convergence rate for row-access methods, Subsection 2.3 provides numerical experiments, and Subsection 2.4 extends TARK to semi-infinite least-squares problems.
@@ 2.1 Proof of main theorem
The proof of Theorem 2 follows the pattern of analysis initiated in [1], but it takes a step further by bounding the inner product terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{t+s}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{t}-\boldsymbol{x}_{\star})\bigr{]}$ which decay exponentially fast with $s$. For ease of reading, the analysis is presented as three lemmas followed by one main calculation.
Lemma 1 (Multi-step expectations).
The RK iteration (2) satisfies
@@
Proof.
For any $t\in\{r,\ldots,s-1\}$, write the one-step update (2b) as
@@
Use the sampling probabilities (2a) to calculate the expectation over the random index $i_{t}$:
@@
Iterating this equation completes the proof. ∎
Lemma 2 (Demmel condition number bound).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration (2) satisfies
@@
Proof.
By the construction of the RK iterates (2b), observe that $\boldsymbol{x}_{r}$ is in the range of $\boldsymbol{A}^{\top}$, as is the solution vector $\boldsymbol{x}_{\star}=\boldsymbol{A}^{+}\boldsymbol{b}=\boldsymbol{A}^{\top}(\boldsymbol{A}\boldsymbol{A}^{\top})^{+}\boldsymbol{b}$. Hence, $\boldsymbol{x}_{r}-\boldsymbol{x}_{\star}\in\operatorname{range}(\boldsymbol{A}^{\top})$. The result follows by expanding $\boldsymbol{x}_{r}-\boldsymbol{x}_{\star}$ in the basis of $\boldsymbol{A}$’s right singular vectors. ∎
Lemma 3 (Mean square errors, based on [1]).
Assume $\boldsymbol{x}_{0}\in\operatorname{range}(\boldsymbol{A}^{\top})$. Then the RK iteration (2) satisfies
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\boldsymbol{x}_{r}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}\leq\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{r}\,\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}+\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$
@@
Proof.
For any $t\in\{0,1,\ldots,r-1\}$, write the one-step update (2b) as
$\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\star}=\underbrace{\biggl{[}\mathbf{I}-\frac{\boldsymbol{a}_{i_{t}}^{\vphantom{\top}}\boldsymbol{a}_{i_{t}}^{\top}}{\lVert\boldsymbol{a}_{i_{t}}\rVert^{2}}\biggr{]}}_{\text{orthogonal projection}}\bigl{(}\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\bigr{)}+\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{\star}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}}.$ (4)
The decomposition explicitly identifies an orthogonal projection matrix. The matrix is idempotent, it annihilates the vector $\boldsymbol{a}_{i_{t}}$, and it preserves all vectors orthogonal to $\boldsymbol{a}_{i_{t}}$. Hence, using the orthogonal decomposition (4) it follows
@@
Use the sampling probabilities (2a) to calculate the expectation over the random index $i_{t}$:
@@
$\leq(1-\kappa_{\rm dem}^{-2})\,\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\rVert^{2}+\frac{\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}},$
where the inequality follows from Lemma 2. Next, take the expectation over the random indices $i_{0},\ldots,i_{t}$:
$\operatorname{\mathbb{E}}\,\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\star}\rVert^{2}\leq(1-\kappa_{\rm dem}^{-2})\,\operatorname{\mathbb{E}}\,\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{\star}\rVert^{2}+\frac{\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}},\quad\text{for each }t\in\{0,\ldots,r-1\}.$
Since $\sum_{s=0}^{\infty}(1-\kappa_{\rm dem}^{-2})^{s}=\kappa_{\rm dem}^{2}=\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\left\|\boldsymbol{A}\right\|_{\rm F}^{2}$, this equation implies the desired result. ∎
Proof of Theorems 2 and 3.
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}=\frac{1}{(t-t_{\rm b})^{2}}\sum\nolimits_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\,\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\star})\bigr{]}.$
Next analyze the terms $\operatorname{\mathbb{E}}\,\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\star})\bigr{]}$ for $r\leq s$ using Lemmas 1, 2, and 3:
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}$ $=\frac{1}{(t-t_{\rm b})^{2}}\sum\nolimits_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\,\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\star})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\star})\bigr{]}$
$\leq\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{t_{\rm b}}\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}+\frac{\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}}{(t-t_{\rm b})^{2}}\sum\nolimits_{r,s=t_{\rm b}}^{t-1}\bigl{(}1-\kappa_{\rm dem}^{-2}\bigr{)}^{\,|s-r|}$
@@
which completes the proof of Theorem 2. Theorem 3 is proved in a similar way, by explicitly averaging over term A also. ∎